# Order Out of Noise and Disorder: Fate of the Frustrated Manifold

### v1

@@ 3.7 Results on the Bounded Cylinder: $S^{1}\times[0,H]\to\mathbb{Z}_{2}$
with no deterministic drift. This Langevin equation predicts pure diffusive dynamics for the cumulative rotation. The theoretical result (68) predicts $\langle(d\Theta_{z})^{2}\rangle=2DR^{2}N\,dt$, giving an effective rotational diffusion coefficient $D_{\text{rot}}=DR^{2}N$. With $N=400$, $D=0.4$, $R=1$, this predicts $D_{\text{rot}}=160$. The strong correlations between vertically aligned clusters ($r>0.8$) indicate that the system behaves approximately like a rigid body, but surprisingly the actual MSD scaling is ballistic ($\alpha\approx 2$) rather than diffusive; a deviation from the theoretical prediction that we attribute to persistent collective modes sustained by the boundary conditions.
@@
Quantitative Rotational Dynamics Analysis.
Figure 4 presents an analysis of rotational dynamics and cluster formation on the bounded cylinder for a simulation with $N=400$ particles at temperature $T=0.4$ with reflective $z$-boundaries, covering the full time range from $t=0$ to $t=50$.
Figure 4: Cylinder: Rotational dynamics analysis for $N=400$ particles, $T=0.4$, Gaussian $\phi_{ij}$, reflective $z$-boundaries, full time series $t=0$ to $50$. (a) Total cumulative rotation $\Theta_{\theta}(t)$ showing large-amplitude fluctuations; vertical dashed line marks cluster formation at $t\approx 1$. (b) Per-cluster cumulative rotation for all four clusters (C0–C3), demonstrating independent fluctuations. (c) $\Theta_{\theta}$ distribution in steady state with $\langle\Theta_{\theta}\rangle=45.8$, $\sigma=243.8$. (d) Angular velocity $\omega_{\theta}$ distribution with $\sigma=0.61$. (e) Rotational energy $E_{\text{rot}}$ time series with mean $\approx 31$. (f) Formation phase detail (t=0t=01515) showing rapid cluster emergence. (g) $\Theta_{\theta}$ autocorrelation with rapid decay ($\tau<0.2$). (h) Inter-cluster angular velocity correlation matrix showing weak anti-correlations ($\rho\approx-0.06$ to 0.21-0.21). (i) Angular MSD with exponent $\alpha\approx 2.0$ (ballistic). (j) Phase comparison: formation (red, $\sigma=228$) vs. steady state (blue, $\sigma=244$).
Cumulative Rotation Statistics.
The total cumulative rotation about the cylinder axis, $\Theta_{\theta}=R^{2}\sum_{i}\dot{\theta}_{i}\,dt$, exhibits large-amplitude fluctuations. Table 8 summarizes the statistics for both the formation and steady-state phases.
Table 8: Cumulative rotation statistics on the cylinder separated by phase (dimensionless units). Unlike the sphere, the standard deviation remains essentially constant across phases, indicating that cluster formation does not significantly constrain the fluctuations.
@@
Full simulation t=0t=05050 $44.6$ $243.8$
The non-zero mean cumulative rotation ($\langle\Theta_{\theta}\rangle\approx 46$) reflects the specific stochastic realization; symmetry only requires $\langle\Theta_{\theta}\rangle=0$ when averaged over initial conditions. The standard deviation $\sigma\approx 244$ is substantially larger than for the sphere ($\sigma\approx 149$). This can be understood from the theoretical framework: the cylinder’s effective diffusion coefficient $D_{\text{rot}}=DR^{2}N$ (Table 1) differs from the sphere’s $D_{\text{rot}}=2ND/3$ due to the different metric structure. Moreover, the four independent clusters contribute additively to the total cumulative rotation variance, amplifying fluctuations compared to the coherent band on the sphere.
Per-Cluster Statistics.
Table 9 shows the cumulative rotation statistics for each of the four clusters in the steady-state phase.
Table 9: Per-cluster cumulative rotation statistics on the cylinder during the steady-state phase (t=1t=15050), in dimensionless units. Each cluster contains approximately 90–100 particles. The similar statistics across clusters reflect the symmetric cluster arrangement.
Cluster Particles $\langle\Theta_{\theta}\rangle$ $\sigma_{\Theta_{\theta}}$
C0 (top) 92 $17.8$ $160.9$
C1 (top) 92 $4.9$ $160.6$
C2 (bottom) 105 $11.9$ $158.5$
C3 (bottom) 96 $15.9$ $157.0$
Inter-Cluster Correlations.
The angular velocity correlation matrix (Figure 4h) reveals the correlation structure between the four clusters:
$\rho_{ij}=\frac{\langle\omega_{i}\omega_{j}\rangle-\langle\omega_{i}\rangle\langle\omega_{j}\rangle}{\sigma_{\omega_{i}}\sigma_{\omega_{j}}}$ (40)
Key findings include:
  • All off-diagonal correlations are weakly negative: $\rho\approx-0.06$ to 0.21-0.21
  • The strongest anti-correlation: C0–C1 ($\rho=-0.21$)
  • These weak anti-correlations are a direct consequence of the theoretical result (68): since $d\Theta_{z}=\sqrt{2D}\cdot R\sum_{i}dW_{i}^{\theta}$ contains only noise terms with no inter-particle coupling through forces, cluster angular velocities are driven by independent noise sources but constrained by the total cumulative rotation evolution, inducing weak anti-correlations
@@
The mean-square displacement (MSD) of the cumulative azimuthal rotation exhibits power-law scaling:
@@
Notably, the cylinder exhibits ballistic behavior ($\alpha\approx 2$), in contrast to the diffusive dynamics ($\alpha\approx 1$) observed on the sphere. This result is robust across multiple independent realizations: ensemble averaging over five runs yields $\alpha=2.00\pm 0.05$, confirming that the ballistic scaling is not a finite-sample artifact but genuine physics.
The ballistic behavior indicates persistent collective motions that do not decorrelate over the simulation timescale. Despite the sample mean $\bar{\omega}$ averaging to approximately zero across realizations (ruling out systematic drift), the four-cluster configuration creates coherent rotational modes that maintain their coherence. This may arise from the boundary conditions: clusters pinned at the reflective $z$-boundaries can only relax collective rotation through azimuthal rearrangement, creating long-lived rotational excitations. The discrete $\mathbb{Z}_{2}$ residual symmetry (two equivalent configurations related by $180^{\circ}$ rotation) may also play a role, as the system cannot continuously explore a Goldstone manifold but instead executes large-amplitude oscillations between the two equivalent states.
@@
The four clusters localize at the reflective boundaries:
  • Clusters C0 and C1: $z\approx+1.5$ (top boundary)
  • Clusters C2 and C3: $z\approx-1.5$ (bottom boundary)
The azimuthal distribution shows that clusters within each level are separated by approximately $90^{\circ}$$180^{\circ}$ in $\theta$, reflecting the azimuthal fragmentation induced by coupling frustration.
@@ 3.8 Results on the Torus: $S^{1}\times S^{1}\to\coprod_{2}S^{1}$
Quantitative Rotational Dynamics Analysis.
Figure 6 presents an analysis of rotational dynamics and ring formation on the torus for a simulation with $N=400$ particles at temperature $T=0.4$, covering the full time range from $t=0$ to $t=50$.
Figure 6: Torus: Rotational dynamics analysis for $N=400$ particles, $T=0.4$, $R=1.5$, $r=0.6$, Gaussian $\phi_{ij}$, full time series $t=0$ to $50$. (a) Total cumulative rotation $\Theta_{\theta}(t)$ showing large-amplitude fluctuations; vertical dashed line marks ring formation at $t\approx 13$. (b) Per-ring cumulative rotation for Ring 0 (red) and Ring 1 (blue). (c) $\Theta_{\theta}$ distribution by phase: formation (red, $\sigma=359$) vs. steady state (blue, $\sigma=396$). (d) Angular velocity $\omega_{\theta}$ distribution with $\sigma=0.46$. (e) Rotational energy $E_{\text{rot}}$ time series with mean $\approx 82$. (f) Formation phase detail (t=0t=01515). (g) $\Theta_{\theta}$ autocorrelation with rapid decay ($\tau<0.2$). (h) Inter-ring $\omega$ correlation over time, with mean $\rho=-0.27$. (i) Angular MSD with exponent $\alpha\approx 1.9$ (superdiffusive). (j) Energy distribution by phase.
Cumulative Rotation Statistics.
The total cumulative rotation $\Theta_{\theta}$ exhibits large-amplitude fluctuations. Table 11 summarizes the statistics for both the formation and steady-state phases.
Table 11: Cumulative rotation statistics on the torus separated by phase (dimensionless units). The standard deviation increases from formation to steady state, indicating that the two-ring structure allows larger collective fluctuations than the dispersed state.
@@
Full simulation t=0t=05050 $74.3$ $387.4$
The torus exhibits the largest rotational fluctuations among the three geometries ($\sigma\approx 396$ vs. $\sigma\approx 149$ for sphere and $\sigma\approx 244$ for cylinder). This hierarchy is consistent with the theoretical predictions of Table 1: the torus has effective diffusion coefficient $D_{\text{rot}}=DN(R^{2}+3r^{2}/2)$ for uniformly distributed particles, which exceeds both the sphere’s $D_{\text{rot}}=2ND/3$ and the cylinder’s $D_{\text{rot}}=DNR^{2}$ for our parameter values ($R=1.5$, $r=0.6$). Additionally, the position-dependent metric factor $(R+r\cos\varphi)^{2}$ in (71) means that particles at different poloidal positions contribute differently to the cumulative rotation, and the two-ring configuration allows the rings to fluctuate somewhat independently, further amplifying the total variance.
Per-Ring Statistics.
The two rings are both active throughout most of the steady-state phase:
  • Ring 0: $\langle\Theta_{\theta}\rangle=15.1$, $\sigma=304.0$, active in 99% of frames
  • Ring 1: $\langle\Theta_{\theta}\rangle=59.7$, $\sigma=340.2$, active in 90% of frames
Inter-Ring Correlation.
The angular velocity correlation between the two rings is:
$\rho(\omega_{\rm Ring\,0},\omega_{\rm Ring\,1})=-0.27$ (42)
This negative correlation indicates that when one ring accelerates, the other tends to decelerate. This behavior follows from the theoretical framework: the total cumulative rotation evolves according to (71) as $d\Theta_{\theta}\approx\sqrt{2D}\sum_{i}(R+r\cos\varphi_{i})\,dW_{i}^{\theta}$, which is a sum over all particles with position-dependent weights. Since each ring occupies a distinct poloidal position $\varphi$, the rings contribute to $\Theta_{\theta}$ with different effective moment arms. The pure-diffusion nature of this evolution (no deterministic drift from forces) means that fluctuations in one ring’s contribution must be partially compensated by the other to maintain consistency with the total $\Theta_{\theta}$ trajectory, inducing anti-correlation. The correlation is stronger than on the cylinder ($\rho\approx-0.1$ to 0.2-0.2) because the torus has only two structures (vs. four clusters), so each structure accounts for a larger fraction of the total cumulative rotation.
@@
The mean-square displacement (MSD) of the cumulative azimuthal rotation exhibits power-law scaling:
@@
The torus exhibits superdiffusive behavior ($\alpha\approx 1.7$), intermediate between the diffusive sphere ($\alpha\approx 1$) and the ballistic cylinder ($\alpha\approx 2$). This result is robust: ensemble averaging over five independent realizations yields $\alpha=1.73\pm 0.10$.
Notably, the torus shows a systematic positive drift: all five realizations exhibit $\bar{\omega}>0$, with ensemble mean $\langle\bar{\omega}\rangle=+0.075$ (significantly different from zero, $p<0.001$). This is in stark contrast to the sphere and cylinder, where $\bar{\omega}$ fluctuates around zero across realizations. The systematic drift indicates that the two-ring configuration on the torus creates a genuine bias toward rotation in one direction.
The drift persists in both the formation phase ($t<13$) and the steady state ($t>13$), ruling out a purely transient formation effect. Analysis of the ring positions reveals several asymmetries: the rings settle at poloidal positions separated by $\sim 170^{\circ}$ rather than exactly $180^{\circ}$, placing one ring on the outer part of the torus (larger effective radius $R+r\cos\varphi$) and one on the inner part (smaller effective radius). The outer ring has both more particles and a larger metric weight, with a combined contribution to $\Theta_{\theta}$ that is $\sim 2.6\times$ larger than the inner ring.
The systematic drift indicates that the two-ring configuration represents a non-equilibrium steady state rather than true thermodynamic equilibrium. While the Stratonovich drift correction (18) ensures detailed balance for free Brownian motion on the torus, the presence of the confining interaction potential and the resulting spatial localization of particles into rings creates a fundamentally different situation. The drift arises from a stochastic ratchet mechanism: particles undergo thermal fluctuations in $\varphi$ around their mean ring positions, and because the effective noise amplitude $\sigma_{\theta}(\varphi)=\sqrt{2D}/(R+r\cos\varphi)$ varies nonlinearly with $\varphi$, these fluctuations are rectified into directed toroidal motion.
Specifically, when a particle fluctuates to smaller $\varphi$ (toward the outer equator), it experiences weaker noise in the $\theta$-direction; when it fluctuates to larger $\varphi$ (toward the inner equator), it experiences stronger noise. This asymmetric coupling between radial fluctuations and azimuthal diffusivity, combined with the asymmetric ring positions (one ring on the outer torus surface at $\varphi\approx 125^{\circ}$, one on the inner surface at $\varphi\approx 295^{\circ}$), breaks the detailed balance condition that would otherwise ensure zero net current. The rings at different poloidal positions experience different effective “noise temperatures” for toroidal motion, and the resulting probability currents do not cancel.
This interpretation is consistent with the theory of Brownian motors and stochastic ratchets [22], where spatially periodic potentials combined with spatially varying noise can rectify thermal fluctuations into directed motion. The fact that all five independent realizations show positive drift (rather than random signs) confirms this is a deterministic consequence of the ring geometry and metric structure, not a statistical fluctuation. The torus geometry thus provides a natural realization of a geometry-induced Brownian motor, where curved metric structure combined with spatially confined particle distributions converts undirected thermal noise into systematic rotation. The contrast between torus and cylinder dynamics is summarized in Table 12.
Table 12: Comparison of collective dynamics on the bounded cylinder versus the torus (dimensionless units). Both geometries exhibit weak anti-correlations between structures, but show qualitatively different MSD scaling: ballistic on the cylinder, superdiffusive on the torus. The torus uniquely exhibits systematic positive drift.
@@
$\langle\Theta_{\theta}\rangle$ (steady) $45.8$ $80.5$
$\sigma_{\Theta_{\theta}}$ (steady) $243.8$ $396.0$
@@
Physical mechanism Boundary pinning + persistent modes Metric asymmetry + ratchet effect
@@
The quality of dimensional reduction can be assessed through the ring widths: Ring 1 has $\sigma_{\theta}=23.8$ px (toroidal), $\sigma_{\varphi}=21.5$ px (poloidal); Ring 2 has $\sigma_{\theta}=26.0$ px, $\sigma_{\varphi}=20.8$ px. The similar widths in both directions indicate rings concentrated on minor circles with residual spread due to thermal fluctuations, consistent with the quasi-1D reduction $T^{2}\to\coprod_{2}S^{1}$.
The torus exhibits position-dependent Gaussian curvature: positive on the outer surface (near $\varphi=0$) and negative on the inner surface (near $\varphi=\pi$). For the parameters $R=1.5$, $r=0.6$ used in our simulations, these effects are moderate and do not strongly bias which minor circles become ring locations.
Additional Ring Analysis.
The torus ring dynamics can be further analyzed through spectral and diffusion methods. Figure 7 presents power spectra and MSD analysis for the two-ring system.
Figure 7: Torus ring dynamics: Spectral and diffusion analysis. Top row: (a) Power spectrum of total $\Theta_{\theta}$ showing characteristic frequency structure. (b) Power spectrum of per-ring $\omega_{\theta}$ with similar spectral shapes for both rings. (c) Cumulative toroidal rotation for each ring (red, green) and their difference $\Delta\theta$ (purple), showing independent drifts. Bottom row: (d) Ring $\omega$ correlation scatter plot confirming $\rho=-0.27$. (e) Relative angle MSD with slope $\approx 1.13$, indicating slightly superdiffusive relative motion. (f) Distribution of relative angular velocity $\Delta\omega_{\theta}$.
@@
The toroidal cumulative rotation $\Theta_{\theta}=\int_{0}^{t}\sum_{i}(R+r\cos\varphi_{i})^{2}\,d\theta_{i}$ dominates the dynamics with $\langle\Theta_{\theta}\rangle=80.5$, $\sigma_{\Theta_{\theta}}=396.0$ in steady state. The large standard deviation indicates strong instantaneous fluctuations. The relative angular velocity between rings has near-zero mean ($\langle\Delta\omega_{\theta}\rangle=-0.016$ rad/t), indicating no systematic relative drift; the rings maintain their diametrically opposite configuration on average. The MSD of the relative angle scales as $\langle\Delta\theta^{2}\rangle\sim t^{1.13}$, indicating approximately diffusive relative motion.

### v2

@@ 3.7 Results on the Bounded Cylinder: $S^{1}\times[0,H]\to\mathbb{Z}_{2}$
with no deterministic drift. For uncorrelated thermal noise and uniform particle distributions, this Langevin equation would predict diffusive MSD scaling ($\alpha=1$). However, the actual MSD exhibits ballistic scaling ($\alpha\approx 2$), as shown below. This deviation from the naive prediction indicates that the boundary-induced cluster configuration creates persistent correlations in the effective thermal torques that are not captured by the simple local Langevin theory. The strong correlations between vertically aligned clusters ($r>0.8$) suggest that the system behaves approximately like a rigid body with long-lived collective rotational modes.
@@
Cumulative Rotation Dynamics.
The total cumulative rotation about the cylinder axis, $\Theta_{\theta}=R^{2}\sum_{i}\dot{\theta}_{i}\,dt$, exhibits large-amplitude fluctuations characteristic of pure noise. Table 8 summarizes the statistics.
Table 8: Cumulative rotation statistics on the cylinder separated by phase (dimensionless units). Unlike the sphere, the standard deviation remains essentially constant across phases.
@@
The standard deviation $\sigma\approx 244$ is substantially larger than for the sphere ($\sigma\approx 149$), reflecting both the different metric structure and the additive contributions from four independent clusters. The autocorrelation decays rapidly ($\tau<0.2$), and inter-cluster angular velocity correlations are weakly negative ($\rho\approx-0.06$ to 0.21-0.21), consistent with the zero-torque property.
@@
Unlike the sphere, the cylinder exhibits ballistic MSD scaling:
@@
Ensemble averaging over five runs confirms $\alpha=2.00\pm 0.05$. This ballistic behavior indicates persistent collective motions that do not decorrelate over the simulation timescale, possibly arising from clusters pinned at the reflective $z$-boundaries or the discrete $\mathbb{Z}_{2}$ residual symmetry preventing continuous exploration of a Goldstone manifold.
@@
The four clusters localize at the reflective boundaries ($z\approx\pm 1.5$), with azimuthal separation of approximately $90^{\circ}$$180^{\circ}$ reflecting coupling frustration.
Cluster Orientation Dynamics.
On the cylinder, the cluster normal $\hat{\mathbf{n}}(t)$ is confined to lie approximately in the $xy$-plane (perpendicular to the cylinder axis), with $\langle n_{z}\rangle\approx 0$ and $\sigma(n_{z})\approx 0.07$. This reflects the geometric constraint: clusters arranged along the $z$-axis produce a distribution whose minimal-extent direction lies in the horizontal plane. Figure 4 shows the orientation dynamics.
  • Rapid reorientation: The correlation time $\tau_{c}\approx 3.2$ is significantly shorter than on the sphere or torus, indicating faster reorientation of the cluster arrangement. This may reflect the discrete nature of the cluster configuration (four clusters) compared to the continuous band on the sphere.
  • Planar precession: The cumulative precession of $\sim 1200^{\circ}$ over the NESS phase corresponds to a rate of $\sim 32^{\circ}$ per unit time, approximately half the sphere’s rate despite the shorter correlation time. The precession is confined to the $xy$-plane rather than exploring the full $S^{2}$.
  • Constant thickness: Unlike the sphere, the cylinder shows minimal change in thickness between formation ($\mathcal{T}\approx 0.36$) and NESS ($\mathcal{T}\approx 0.34$), suggesting that the cluster configuration establishes quickly and maintains a consistent geometry.
  • Power spectrum: The structure normal power spectrum $S_{n_{z}}(\omega)\propto\omega^{-2.0}$ is the only geometry approaching the theoretical $1/\omega^{2}$ prediction for type-A Goldstone diffusion modes.
Figure 4: Cylinder $S^{1}\times[0,H]$: Cluster orientation dynamics for $N=400$ particles, $T=0.4$, full time series $t=0$ to $50$. Top row: Cluster normal components showing confinement to the $xy$-plane ($n_{z}\approx 0$). Second row: Thickness remains approximately constant ($\mathcal{T}\approx 0.34$); polar angle $\theta\approx 90^{\circ}$ confirming horizontal orientation; cumulative precession at $\sim 32^{\circ}$/time. Third row: Orientation autocorrelation with short correlation time $\tau_{c}\approx 3.2$; component autocorrelations; thickness autocorrelation. Bottom row: Power spectra with $S_{n_{z}}(\omega)\propto\omega^{-2.0}$, the only geometry approaching the theoretical $1/\omega^{2}$ prediction; 3D trajectory confined to the equatorial band of $S^{2}$.
Summary: Cylinder Dynamics.
On the bounded cylinder, dimensional reduction produces four localized clusters (two vertical pairs) that form rapidly by $t\approx 1$. The dynamics differ qualitatively from the sphere: the cumulative rotation exhibits ballistic scaling ($\alpha=2.0$) rather than diffusive, and the orientation correlation time is short ($\tau_{c}\approx 3.2$) compared to the sphere ($\tau_{c}\approx 18$). The ballistic MSD suggests persistent collective modes that do not decorrelate, while the rapid orientation reorientation reflects the discrete nature of the cluster configuration. The boundaries play a key role, pinning clusters at $z\approx\pm 1.5$ and creating the discrete $\mathbb{Z}_{2}$ residual symmetry.
@@ 3.8 Results on the Torus: $S^{1}\times S^{1}\to\coprod_{2}S^{1}$
Cumulative Rotation Dynamics.
Table 10 summarizes the cumulative rotation statistics for the torus.
Table 10: Cumulative rotation statistics on the torus separated by phase (dimensionless units). The standard deviation increases from formation to steady state, indicating that the two-ring structure allows larger collective fluctuations.
@@
The torus exhibits the largest rotational fluctuations among the three geometries ($\sigma\approx 396$ vs. $\sigma\approx 149$ for sphere and $\sigma\approx 244$ for cylinder), consistent with the theoretical predictions of Table 1. The inter-ring angular velocity correlation is $\rho=-0.27$, indicating that when one ring accelerates, the other tends to decelerate. This negative correlation is stronger than on the cylinder ($\rho\approx-0.1$ to 0.2-0.2) because the torus has only two structures versus four clusters.
@@
The MSD of the cumulative azimuthal rotation exhibits power-law scaling:
@@
The torus exhibits superdiffusive behavior, intermediate between the diffusive sphere ($\alpha\approx 1$) and the ballistic cylinder ($\alpha\approx 2$). Ensemble averaging over five independent realizations yields $\alpha=1.73\pm 0.10$.
The torus also shows a systematic positive drift: all five realizations exhibit $\bar{\omega}>0$, with ensemble mean $\langle\bar{\omega}\rangle=+0.075$ (significantly different from zero, $p<0.001$). This is in stark contrast to the sphere and cylinder, where $\bar{\omega}$ fluctuates around zero. The drift arises from a stochastic ratchet mechanism: particles undergo thermal fluctuations in $\varphi$ around their mean ring positions, and because the effective noise amplitude $\sigma_{\theta}(\varphi)=\sqrt{2D}/(R+r\cos\varphi)$ varies nonlinearly with $\varphi$, these fluctuations are rectified into directed toroidal motion. The asymmetric ring positions (one ring on the outer torus surface, one on the inner surface) break detailed balance, creating a geometry-induced Brownian motor [26].
The contrast between torus and cylinder dynamics is summarized in Table 11.
Table 11: Comparison of collective dynamics on the bounded cylinder versus the torus (dimensionless units). Both geometries exhibit weak anti-correlations between structures, but show qualitatively different MSD scaling.
@@
The quality of dimensional reduction can be assessed through the ring widths, which show similar spread in both toroidal and poloidal directions ($\sigma\approx 20$$26$ pixels), consistent with the quasi-1D reduction $T^{2}\to\coprod_{2}S^{1}$.
@@
Ring Orientation Dynamics.
On the torus, the ring normal exhibits qualitatively different behavior: a strong bias toward the $z$-axis with $\langle n_{z}\rangle\approx 0.73$ and large fluctuations $\sigma(n_{z})\approx 0.44$. This reflects the torus geometry: when particles form rings at constant poloidal angle $\varphi$, the plane containing these rings is tilted relative to the $xy$-plane, producing a ring normal with significant $z$-component. Figure 7 illustrates this behavior.
  • Geometric constraint: The mean $\langle n_{z}\rangle\approx 0.73$ corresponds to a ring-plane tilt of $\arccos(0.73)\approx 43^{\circ}$ from horizontal. This is consistent with rings forming at poloidal positions $\varphi\approx\pm 45^{\circ}$ from the outer equator of the torus.
  • Large orientation fluctuations: The standard deviation $\sigma(n_{z})\approx 0.44$ is substantially larger than for the sphere ($\sigma(n_{z})\approx 0.11$), indicating that the ring plane can tilt significantly as particles redistribute between the two rings.
  • Long correlation time: Despite the large fluctuations, the correlation time $\tau_{c}\approx 18$ matches the sphere, suggesting similar underlying dynamics for the slow orientation evolution.
  • Rapid precession: The cumulative precession of $\sim 2400^{\circ}$ (rate $\sim 65^{\circ}$/time) slightly exceeds the sphere, consistent with the systematic drift observed in the cumulative rotation.
  • Power spectrum: The structure normal power spectrum $S_{n_{z}}(\omega)\propto\omega^{-1.4}$ is the shallowest among the three geometries, deviating most from the theoretical $1/\omega^{2}$ prediction for type-A Goldstone diffusion modes.
Figure 7: Torus $T^{2}$: Ring orientation dynamics for $N=400$ particles, $T=0.4$, full time series $t=0$ to $50$. Top row: Ring normal components showing strong $z$-bias ($\langle n_{z}\rangle\approx 0.73$) with large fluctuations. Second row: Thickness remains constant ($\mathcal{T}\approx 0.27$); polar angle $\theta\approx 43^{\circ}$ reflecting the tilted ring plane; cumulative precession at $\sim 65^{\circ}$/time. Third row: Orientation autocorrelation with correlation time $\tau_{c}\approx 18$, comparable to the sphere; component autocorrelations showing different decay rates; thickness autocorrelation. Bottom row: Power spectra with $S_{n_{z}}(\omega)\propto\omega^{-1.4}$ scaling; 3D trajectory on $S^{2}$ showing exploration biased toward the northern hemisphere.
Summary: Torus Dynamics.
On the torus, dimensional reduction produces two rings at diametrically opposite poloidal positions, forming by $t\approx 13$ (the slowest among the three geometries). The cumulative rotation exhibits superdiffusive scaling ($\alpha=1.7$) and systematic positive drift ($\langle\bar{\omega}\rangle=+0.075$), arising from a stochastic ratchet mechanism created by the non-uniform metric. The ring normal shows a strong $z$-bias ($\langle n_{z}\rangle\approx 0.73$) reflecting the tilted ring-plane geometry, with long correlation time ($\tau_{c}\approx 18$) comparable to the sphere. The torus provides the clearest example of how non-uniform local geometry can induce non-equilibrium steady states with directed motion from thermal fluctuations.